# Fundamental Limits of Quantum Semantic Communication via Sheaf Cohomology

### v1

@@ I Introduction
Semantic communication (SC) has emerged as a paradigm shift in communication theory, moving beyond Shannon’s classical framework of transmitting bits to transmitting meaning [9]. This approach promises bandwidth savings in multi-agent systems where users share common objectives but may maintain heterogeneous internal representations shaped by distinct sensing modalities and AI architectures. However, this heterogeneity poses a fundamental challenge: when users interpret symbols through different semantic lenses, perfect bit-level transmission no longer guarantees mutual understanding.
Consider two robots coordinating a manipulation task. Each robot maintains an internal world model encoding spatial relationships, object properties, and action affordances. When one robot communicates “move left,” the meaning depends on reference frames, learned associations, and task context that may differ between agents. The robots face an irreducible ambiguity stemming not from channel noise, but from their heterogeneous representations. This motivates our central questions: What is the minimum communication rate to resolve such ambiguity? Can shared quantum resources help? Classical SC models meanings as vectors in real spaces, with alignment achieved when projections agree [15]. But this framework cannot capture contextuality: the dependence of meaning on the broader context of interpretation. Abramsky and Brandenburger [16] showed that quantum contextuality corresponds precisely to obstructions in sheaf cohomology. By extending semantic spaces to complex Hilbert spaces and semantic channels to quantum operations [13], we gain access to this rich structure. The payoff is twofold: entanglement provides provable advantages for semantic alignment, and cohomological tools characterize when alignment is fundamentally impossible.
@@ I-A Related Work
Despite progress in deep learning for semantic compression [21] and [1], the information-theoretic foundations of SC remain incomplete. Semantic information theory, from Carnap and Bar-Hillel [3] to Floridi [10], addresses what semantic content is. A Recent work [11] generalizes Shannon’s rate-distortion theory by replacing distortion constraints with semantic constraints via truth functions. But these works [10, 11] do not address how to align heterogeneous semantic interpretations across users. The missing piece is a mathematics for semantic consistency that characterizes when locally coherent agents can achieve global alignment, and quantifying obstructions when they cannot. Sheaf theory, recently applied to multi-agent networks [8, 4, 6], formalizes semantic consistency through global sections representing alignment and cohomology measuring irreducible ambiguity. The sheaf-theoretic approach to quantum contextuality [16] was extended cohomologically by [17] and [5]. Entanglement-assisted classical capacity was established by [2]; quantum SC was proposed by [12]. Unlike prior work that assumes shared semantics or proposes SC protocols without capacity theorems, we establish fundamental limits: sheaf cohomology characterizes irreducible semantic ambiguity, and entanglement provably reduces this obstruction.
@@ I-B Our Contributions
The main contribution is a rigorous framework proving that sheaf cohomology precisely characterizes irreducible semantic ambiguity, and that shared entanglement reduces the communication rate required to align heterogeneous agents. Our key contributions include:
  • Cohomological semantic rate (Theorem 1): We prove that the minimum communication rate for perfect semantic alignment equals $\dim H^{1}$, the dimension of the first sheaf cohomology group. This captures irreducible semantic ambiguity and provides a semantic analog of Shannon’s channel coding theorem.
  • Entanglement-assisted capacity (Theorem 2): We derive semantic channel capacity when sender and receiver share prior entanglement, proving it strictly exceeds classical capacity. The quantum advantage grows as channel noise increases—precisely when semantic communication most benefits over bit-level transmission.
  • Contextuality as resource (Theorem 3): We show that quantum contextuality reduces cohomological obstructions to semantic alignment. Contextual correlations act as “pre-shared semantic resolution,” establishing contextuality as a resource for semantic communication.
  • Discord-integration duality (Theorem 4): We prove quantum discord equals integrated semantic information, linking quantum correlations to irreducible semantic content and connecting our framework to integrated information theory.
@@ II Modeling Semantic Networks as Quantum Sheaves
This section establishes a mathematical framework for quantum SC by extending network sheaves from classical vector spaces to quantum Hilbert spaces, thereby enabling the use of tools from algebraic topology and quantum information theory.
@@ II-B Quantum Semantic Sheaves
    $\mathcal{F}_{e}:\mathcal{L}(\mathcal{H}_{u})\to\mathcal{L}(\mathcal{H}_{v})$ is a completely positive trace-preserving (CPTP) map for each edge $e=(u,v)$, representing the semantic channel from agent $u$ to agent $v$.
@@
    $\rho_{v}$ is a density operator on $\mathcal{H}_{v}$ representing agent $v$’s current semantic state.
@@
A local section over $U\subseteq V$ is a semantic configuration $\sigma$ such that $\mathcal{F}_{e}(\sigma_{u})=\sigma_{v}$, for all edges $e=(u,v)$ with both endpoints in $U$. A global section is a local section over the entire vertex set $V$.
When a global section exists, the network achieves perfect semantic alignment. This means that all agents share a consistent semantic state that transforms coherently along every communication link. The central question we address here is: when does such alignment exist, and what communication resources are required to achieve it?
@@ II-C The Sheaf Laplacian
$\mathcal{L}=(\delta^{0})^{\dagger}\delta^{0},$
where $\dagger$ denotes the adjoint with respect to the Hilbert–Schmidt inner product. The kernel of $\mathcal{L}$ consists precisely of the global sections.
@@ II-D Cohomological Obstruction
To formally characterize irreducible semantic obstructions, we construct the Čech complex adapted to our quantum setting:
$C^{0}(G,S)\xrightarrow{\ \delta^{0}\ }C^{1}(G,S)\xrightarrow{\ \delta^{1}\ }C^{2}(G,S)\xrightarrow{\ \delta^{2}\ }\cdots,$ (2)
where $C^{0}$ consists of vertex assignments (semantic configurations), $C^{1}$ consists of edge assignments, and the coboundary maps $\delta^{k}$ measure inconsistency at each level. Cohomology groups are defined as the quotient of cycles (elements in the kernel) by boundaries (elements in the image), capturing obstructions that cannot be removed by local adjustments. For semantic alignment:
@@
$H^{1}(G,\mathcal{S})$ $=\ker(\delta^{1})/\mathrm{im}(\delta^{0})$ $\text{(obstruction space)}.$ (4)
The zeroth cohomology $H^{0}$ captures what global alignments exist. The first cohomology $H^{1}$ captures irreducible obstructions, which are semantic ambiguities that cannot be resolved by any local processing. A nonzero element of $H^{1}$ represents a configuration that is locally consistent on every edge but cannot be extended to a global section.
@@ II-F Connection to Contextuality
Our framework connects directly to quantum contextuality [16]. In their formulation, a measurement scenario defines a sheaf over a simplicial complex of compatible measurements, and contextuality corresponds to the nonexistence of global sections.
@@ III-B Entanglement-Assisted Semantic Capacity
$\dim H^{1}_{EA}(G,\mathcal{S})=\max\left(0,\dim H^{1}(G,\mathcal{S})-\sum_{e\in E}\log_{2}r_{e}\right).$ (7)
@@
These results establish a precise connection between algebraic topology and communication theory. The key distinction is that Shannon theory concerns statistical uncertainty (entropy), while our framework concerns structural uncertainty (cohomology) which are obstructions arising from the topology of heterogeneous representations rather than probabilistic noise.
@@ IV-A Contextuality as Semantic Resource
IV-B Quantifying Contextual Advantage
$\mathrm{CF}(e)=\min_{e^{\prime}\in\mathrm{NC}}\|e-e^{\prime}\|_{1}$ (9)
@@ IV-B Quantifying Contextual Advantage
The semantic rate reduction due to contextuality satisfies: $R_{\text{classical}}-R_{\text{quantum}}\geq CF(e)\textperiodcentered\log_{2}\,\text{dim}\,H^{1}_{\text{classical}}$
@@ IV-C Quantum Discord and Semantic Integration
Both quantities capture the same irreducibility. The quantum mutual information decomposes as $I(A:B)=J(A:B)+\delta(A:B)$. The classical correlation $J(A:B)$ corresponds to the partitioned information $\sum_{i}I(A_{i}:B_{i})$ under the optimal measurement-induced partition. The discord $\delta(A:B)$ is what remains: information that cannot be captured by any local measurement strategy. Conversely, the integrated information $\Phi_{\mathrm{sem}}$ seeks the minimum information partition (MIP), which for quantum states corresponds to the optimal measurement basis. The “leftover” after optimal classical extraction is exactly the quantum discord. ∎
@@ V Conclusion
We developed an information-theoretic framework for quantum semantic communication using sheaf cohomology. Our main results show that: (1) the minimum semantic alignment rate equals $\log_{2}\dim H^{1}$, (2) entanglement reduces this rate by providing a rigorous mechanism for shared context, (3) contextuality further reduces communication requirements, and (4) quantum discord quantifies irreducible semantic content. Future work includes experimental validation, extension to continuous-variable systems, and application to multi-agent AI alignment.

### v2

@@ I Introduction
Semantic communication (SC) has emerged as a paradigm shift in communication theory, moving beyond Shannon’s classical framework of transmitting bits to transmitting meaning [12]. This approach promises bandwidth savings in multi-agent systems where users share common objectives but may maintain heterogeneous internal representations shaped by distinct sensing modalities and artificial intelligence (AI) architectures. However, this heterogeneity poses a fundamental challenge: when users interpret symbols through different semantic lenses, perfect bit-level transmission no longer guarantees mutual understanding.
Consider two robots coordinating a manipulation task. Each robot maintains an internal world model encoding spatial relationships, object properties, and action affordances. When one robot communicates “move left,” the meaning depends on reference frames, learned associations, and task context that may differ between agents. The robots face an irreducible ambiguity stemming not from channel noise, but from their heterogeneous representations. This motivates our central questions: What is the minimum communication rate to resolve such ambiguity? Can shared quantum resources help? Classical SC models meanings as vectors in real spaces, with alignment achieved when agents’ representations, projected onto a shared semantic space, agree [18]. But this framework cannot capture contextuality, where a symbol’s meaning depends on what other symbols are jointly queried and the history of prior exchanges. Static vector projections [18] cannot represent such dependencies. Mathematically, contextuality manifests as a topological obstruction: local consistencies that cannot be extended globally. Abramsky and Brandenburger [19] formalized this insight, showing that quantum contextuality corresponds precisely to obstructions in sheaf cohomology. By extending semantic spaces to complex Hilbert spaces and semantic channels to quantum operations [16], we gain access to this rich structure. The resulting quantum framework provides both a resource for alignment and a diagnostic tool: entanglement reduces required communication, while sheaf cohomology characterizes irreducible obstructions.
@@ I-A Related Work
Despite progress in deep learning for semantic compression [1, 27, 8, 26], the information-theoretic foundations of SC remain incomplete. Semantic information theory, from Carnap and Bar-Hillel [3] to Floridi [13], addresses what semantic content is. A recent work [14] generalizes Shannon’s rate-distortion theory by replacing distortion constraints with semantic constraints via truth functions. But these works [13, 14] do not address how to align heterogeneous semantic interpretations across users. The missing piece is a mathematics for semantic consistency that characterizes when locally coherent agents can achieve global alignment, and quantifying obstructions when they cannot. Sheaf theory, recently applied to multi-agent networks [11, 7, 6], formalizes semantic consistency through global sections representing alignment and cohomology measuring irreducible ambiguity. The sheaf-theoretic approach to quantum contextuality [19] was further developed by [20] and [5] using sheaf cohomology, an algebraic topological tool that rigorously measures obstructions to extending local consistency to global agreement. Entanglement-assisted classical capacity was established by [2]; quantum SC was proposed by [15]. Unlike prior work that assumes shared semantics or proposes SC protocols without capacity theorems, we establish fundamental limits.
@@ I-B Our Contributions
The main contribution of this paper is a rigorous framework proving that sheaf cohomology precisely characterizes irreducible semantic ambiguity, and that shared entanglement reduces the communication rate required to align heterogeneous agents. Our key contributions include:
  • Cohomological semantic rate (Theorem 1): We prove that the minimum communication rate for perfect semantic alignment equals $\dim H^{1}$, the dimension of the first sheaf cohomology group. Intuitively, $\dim H^{1}$ counts the number of independent semantic ambiguities that cannot be resolved by local processing alone. This provides a semantic analog of Shannon’s channel coding theorem [21].
  • Entanglement-assisted capacity (Theorem 2): We derive semantic channel capacity when sender and receiver share prior entanglement, proving it strictly exceeds classical capacity. Entanglement provides a physical realization of the “shared context” often assumed in classical SC, with each shared entangled pair saving one bit of alignment communication. The quantum advantage grows as channel noise increases, which is precisely when SC most benefits over bit-level transmission [4].
  • Contextuality as resource (Theorem 3): We show that quantum contextuality reduces cohomological obstructions to semantic alignment. Contextual correlations act as “pre-shared semantic resolution,” establishing contextuality as a resource for SC.
  • Discord-integration duality (Theorem 4): We prove that quantum discord, which measures quantum correlations beyond entanglement, equals integrated semantic information. This links quantum correlations to irreducible semantic content and connects our framework to integrated information theory [25].
@@ II Modeling Semantic Networks as Quantum Sheaves
This section establishes a mathematical framework for quantum SC by extending network sheaves from classical vector spaces to quantum Hilbert spaces, thereby enabling the use of tools from algebraic topology and quantum information theory [16].
@@ II-B Quantum Semantic Sheaves
    $\mathcal{F}_{e}:\mathcal{L}(\mathcal{H}_{u})\to\mathcal{L}(\mathcal{H}_{v})$ is a completely positive trace-preserving (CPTP) map for each edge e=(u,v)e\!=\!(u,v), representing the semantic channel from agent $u$ to $v$.
@@
    $\rho_{v}$ is a density operator on $\mathcal{H}_{v}$ representing agent $v$’s current semantic state. A density operator is a positive semi-definite matrix with $\mathrm{Tr}(\rho)=1$ , generalizing probability distributions to quantum systems [16].
@@
A local section over a subset of agents $U\subseteq V$ is a semantic configuration $\sigma$ such that $\mathcal{F}_{e}(\sigma_{u})=\sigma_{v}$, for all edges $e=(u,v)$ with both endpoints in $U$. A global section is a local section over the entire vertex set $V$.
When a global section exists, the network achieves perfect semantic alignment. This means that all agents share a consistent semantic state that transforms coherently along every communication link. The central question we address here is: when does such alignment exist, and what communication resources are required to achieve it? To answer this, we first need tools to measure how far a configuration is from perfect alignment, and to characterize obstructions when perfect alignment is impossible.
@@ II-C The Sheaf Laplacian
$\mathcal{L}=(\delta^{0})^{\dagger}\delta^{0},\vskip-2.84526pt$ (2)
where $\dagger$ denotes the adjoint operator with respect to the Hilbert–Schmidt inner product $\langle A,B\rangle_{\text{HS}}=\text{Tr}(A^{\dagger}B)$. The adjoint $(\delta^{0})^{\dagger}$ is the unique linear operator satisfying the duality condition: $\langle(\delta^{0})^{\dagger}x,y\rangle_{\text{HS}}=\langle x,\delta^{0}y\rangle_{\text{HS}},$ for all semantic configurations $x,y$. The operator $\delta^{0}$ maps vertex assignments to edge disagreements; $(\delta^{0})^{\dagger}$ reverses this flow by aggregating edge errors back to vertices. The kernel of $\mathcal{L}$ consists precisely of the global sections:
$\ker(\mathcal{L})=\{\sigma:F_{e}(\sigma_{u})=\sigma_{v}\text{ for all edges }e=(u,v)\}.\vskip-2.84526pt$ (3)
@@ II-D Cohomological Obstruction
To formally characterize irreducible semantic obstructions, we construct the Čech complex [9] adapted to our quantum setting:
$C^{0}(G,S)\xrightarrow{\ \delta^{0}\ }C^{1}(G,S)\xrightarrow{\ \delta^{1}\ }C^{2}(G,S)\xrightarrow{\ \delta^{2}\ }\cdots,\vskip-2.84526pt$ (4)
where $C^{0}$ consists of vertex assignments that are semantic states, $C^{1}$ consists of edge assignments that are semantic disagreements along communication links, and the coboundary maps $\delta^{k}$ measure inconsistency at each hierarchical level (shared semantics among $k+1$ set of agents). Cohomology groups are further defined as the quotient of cycles (elements in the kernel) by boundaries (elements in the image), capturing obstructions that cannot be removed by local adjustments. For semantic alignment:
@@
$H^{1}(G,\mathcal{S})$ $=\ker(\delta^{1})/\mathrm{im}(\delta^{0})$ $\text{(obstruction space)}.\vskip-5.69054pt$ (6)
The zeroth cohomology $H^{0}$ captures what global alignments exist. The first cohomology $H^{1}$ captures irreducible obstructions, which are semantic ambiguities that cannot be resolved by any local processing. A nonzero element of $H^{1}$ represents a configuration that is locally consistent on every edge but cannot be extended to a global section. To illustrate this, consider three agents $A,B,C$ in a loop, each maintaining different semantic reference frames. A 1-cocycle assigns frame disagreements to edges: $(d_{AB},d_{BC},d_{CA})$. If these satisfy $d_{AB}+d_{BC}+d_{CA}=0$ (loop consistency), they form a cycle. However, if no global choice of reference frames can simultaneously satisfy all edges, then this cycle represents a topological incompatibility, i.e., an element of $H^{1}\neq 0$. This obstruction cannot be removed by local communication; it reflects fundamental misalignment in the network structure.
@@ II-F Connection to Contextuality
Our framework connects directly to quantum contextuality [19]. In [19], a measurement scenario defines a sheaf over a simplicial complex of compatible measurements, and contextuality corresponds to the nonexistence of global sections.
@@ III-B Entanglement-Assisted Semantic Capacity
$\dim H^{1}_{EA}(G,\mathcal{S})=\max\left(0,\dim H^{1}(G,\mathcal{S})-\sum_{e\in E}\log_{2}r_{e}\right).\vskip-2.84526pt$ (10)
@@
These results establish a precise connection between algebraic topology and information theory. The key distinction is that Shannon theory concerns statistical uncertainty (entropy), while our framework concerns structural uncertainty (cohomology) which are obstructions arising from the topology of heterogeneous representations rather than probabilistic noise.
@@ IV-A Contextuality as Semantic Resource
IV-B Quantifying Contextual Advantage
$\mathrm{CF}(e)=\min_{e^{\prime}\in\mathrm{NC}}\|e-e^{\prime}\|_{1}\vskip-2.84526pt$ (11)
@@ IV-B Quantifying Contextual Advantage
The semantic rate reduction due to contextuality satisfies: $R_{\text{classical}}-R_{\text{quantum}}\geq CF(e)\textperiodcentered\log_{2}\,\text{dim}\,H^{1}_{\text{classical}}$.
@@ IV-C Quantum Discord and Semantic Integration
We show the equality by demonstrating that both quantities measure the same irreducible correlations from dual perspectives. The quantum mutual information decomposes as:
$I(A:B)=J(A:B)+\delta(A:B),$ (15)
where $J(A:B)=\max_{M_{B}}[S(\rho_{A})-S(\rho_{A}|M_{B})]$ is the classical correlation extractable by measuring subsystem B, and $\delta(A:B)$ is the quantum discord. The classical correlation $J(A:B)$ represents the maximum information about A obtainable through local measurements on B. This corresponds precisely to $\max_{\mathrm{partitions}}\sum_{i}I(A_{i}:B_{i})$ in the definition of $\Phi_{\mathrm{sem}}$, since the optimal measurement on B induces the partition that maximizes extractable local information. Substituting into the definition of $\Phi_{\mathrm{sem}}$:
$=I(A:B)-J(A:B)=\delta(A:B).$
Thus both quantities answer the same question from dual perspectives: discord asks “what correlations survive any local measurement?” while integrated information asks “what information cannot be decomposed into parts?” The equality holds because the optimal measurement basis for extracting classical correlations coincides with the minimum information partition. ∎
@@ V Conclusion
We developed an information-theoretic framework for quantum SC using sheaf cohomology. Our main results show that: (1) the minimum semantic alignment rate equals $\log_{2}\dim H^{1}$, (2) entanglement reduces this rate by providing a rigorous mechanism for shared context, (3) contextuality further reduces communication requirements, and (4) quantum discord quantifies irreducible semantic content. Future work includes experimental validation, extension to continuous-variable systems, and application to multi-agent AI alignment.